# A Symmetric Multigrid-Preconditioned Krylov Subspace Solver for Stokes Equation

### v1

@@ 3.2 Symmetric Vanka smoother
The Vanka smoother is more powerful but with a higher computational cost per iteration compared to the distributive smoother. It can be viewed as a specific type of Schwarz smoother schoberl2003schwarz . The fundamental idea behind the Vanka smoother is to solve the Stokes equations locally, subdomain by subdomain, wherein all degrees of freedom within one subdomain are updated simultaneously. Similar to the other smoother, the Vanka smoother can be expressed in the format of Eq. 3. In the Vanka smoother, we divide the entire problem domain $D$ into $n$ potentially overlapping subdomains $\{D_{1},...,D_{n}\}$ as shown in Fig. 4. Each subdomain can be viewed as a local saddle-point problem. When solving each subdomain problem sequentially, akin to Gauss-Seidel, it’s called the multiplicative Vanka smoother. The smoother operator can be defined as
@@
where $\mathbf{C}_{i}$ represents the selection matrix from domain $D$ to domain $D_{i}$. Alternatively, if each subdomain problem is solved simultaneously as Jacobi method, it’s the addictive Vanka smoother. The smoother operator for this approach is defined as
@@
We prove that the addictive Vanka smoother is always symmetric in B. To ensure the symmetry in the multiplicative Vanka smoother, we adopt a similar approach to symmetric Gauss-Seidel. We solve each subdomain from domain $D_{1}$ to $D_{n}$, and then in reverse order from $D_{n}$ to $D_{1}$ again. The symmetry for our symmetric multiplicative Vanka smoother is shown in C.
@@ 3.3 Symmetric integration of smoother
As previously discussed, these two smoother provide advantages depending on the region of the computational domain. Symmetric distributive smoother demonstrates reliability when applied to the interior region, while symmetric Vanka smoother exhibits greater suitability near the boundary. However, despite each can be made symmetric on its own, it’s not universally true that any combination will inherently retain symmetry. To ensure strict symmetry with the multigrid framework, we propose the following smoothing step algorithm. First, we partition the unknowns into two non-overlapping sets: boundary set $V_{1}$ and interior set $V_{2}$. Then we apply symmetric Vanka smoother to $V_{1}$, symmetric distributive smoother to $V_{2}$ and another symmetric Vanka smoother to $V_{1}$ again. The detailed algorithm is shown in Alg. 3, with a formal proof for its symmetry provided in D. Practically, DGS smoother is used instead of DWJ smoother because it has a faster convergence rate but a similar time cost.
This approach leverages the strengths of both smoothers. On one hand, the larger size of $V_{2}$ ensures that symmetric Vanka smoother isn’t overused, enhancing computational efficiency. On the other hand, employing symmetric Vanka smoother near the boundary helps prevent the potential divergence issues of symmetric distributive smoother. Therefore, this integration approach is a superior choice in practical applications compared to using either smoother alone.
@@
3:procedure Smoothing($\mathbf{L}^{h}$, $\mathbf{x}^{h}_{0}$, $\mathbf{b}^{h}$)
@@ 4 Preconditioning
Although the use of Vanka smoother significantly improves convergence near the boundary, multigrid is more stable to be used as the preconditioner for the Krylov subspace solver. SQMR is known for its lower computational time and storage requirements, and supports arbitrary symmetric preconditioners, including indefinite ones from multigrid V-cycle. Therefore, we propose using our multigrid as the preconditioner for SQMR. The multigrid preconditioner is outlined in Alg. 4. It’s worth mentioning that not any multigrid V-cycle can be used as the preconditioner. This procedure must be equivalent to $\mathbf{t}\leftarrow\mathbf{M}^{-1}\mathbf{s}_{0}$ with some symmetric matrix $\mathbf{M}$. Fortunately, as we prove in the appendix, our multigrid indeed satisfies this condition with some initial guess $\mathbf{x}_{0}=\mathbf{0}$.
@@
where $\gamma$ is a small enough number. This modification is corresponding to the classical penalty method. Since the modification is only applied as a preconditioner instead of the system matrix. The divergence-free property can be always exactly maintained.
@@
Algorithm 4 Multigrid-Preconditioned SQMR
@@ 5 Discrete Domain Design
In this section, we describe our discrete domain designed for boundary settings. Every cell within the domain is labelled as Dirichlet, exterior or interior as illustrated in Fig. 5. We assign Dirichlet boundary conditions to velocity variables located on the surface of Dirichlet cells. Velocity variables on the surfaces of at least one interior cell, and not on any Dirichlet cell, are considered active degrees of freedom and are included in our discretization process. Any other velocity variable is not a true degree of freedom and is excluded from our discretization. Pressure variables are treated as degrees of freedom if they are within an interior cell; otherwise, they are not. We adopt this cell-based designation, as opposed to, e.g. designating individual grid faces as Neumann faces, as this design leads itself naturally to coarsening this cell-level designation as follows: A cell at a coarser level inherits its label based on the labels of the finer-level cells within it. Specifically, a coarser-level cell is labelled as Dirichlet if it encloses any finer-level Dirichlet cell, otherwise as interior if it contains any finer-level interior cell. It’s still an exterior cell if all enclosed finer-level cells are exterior. This discrete domain design enables a straightforward partition of degrees of freedom into two non-overlapping sets for the use of symmetric Vanka and symmetric distributive smoothers. If a cell is proximal to any Dirichlet cell or exterior cell, all the degrees of freedom on this cell are smoothed by symmetric Vanka smoother. Conversely, symmetric distributive smoother is used for all other degrees of freedom.
Lastly, we address the formulation of equations for each degree of freedom. We apply a finite-difference method to discretize Stokes equations for active velocity variables between two interior cells and active pressure variables. Velocity variables between an interior and a Dirichlet cell are governed by Dirichlet boundary conditions. For velocity variables between an interior and an exterior cell, we assume Neumann boundary conditions for velocity gradients and consider the missing pressures as zero.
@@
6 Numerical experiments
In this section, we present the results of our numerical experiments, which compare the performance of our method with the pure multigrid method under different boundary conditions. The fluid viscosity is set to $\eta=10^{-3}m^{2}/s$, and $\gamma=10^{-3}$. To enhance the efficiency of Vanka smoothing, we pre-factorize $3^{4}$ in 2D or $3^{6}$ in 3D local saddle point problems, based on whether neighboring cells are Dirichlet, exterior or interior. It’s direct to parallel addictive Vanka smoother, and an 8-color scheme is applied to parallel multiplicative Vanka smoother. All examples are implemented in C++ code, and we utilize Intel Pardiso and Eigen as the direct solvers for the bottom level of V-cycle and for solving local saddle-point problems in the Vanka smoother respectively. The experiments are performed on a computer with an Intel(R) Xeon(R) CPU E5-2650 v3 @ 2.30GHz and 128 GB of main memory. Paraview is used for postprocessing the results. In the following section, we’ll discuss the results in 2D examples and 3D examples separately.
@@ 6 Numerical experiments
In this subsection, we explore four 2D examples to assess the performance of our approach. Specifically, we examine the driven cavity example and Poiseuille flow around cylinder example to evaluate the convergence rates in comparison to pure multigrid methods. Additionally, the remaining two examples demonstrate the capability of our method to effectively handle more complex scenarios beyond the scope of pure multigrid methods.
@@ 6.1.1 Driven Cavity Example
The driven-cavity example is a well-established benchmark in the field of computational fluid. In the 2D scenario, the computational domain is a unit square, where no-slip boundary conditions are imposed on the left, right and bottom sides. While on the top side, a zero vertical velocity and a constant horizontal velocity of $u=1.0m/s$ are imposed. We use a $1024^{2}$ unit square as our computational domain and an 8-level multigrid as our preconditioner and solver. The solved velocity field with streamline and the convergence of pure multigrid on $\mathbf{\tilde{L}}$ and multigrid-preconditioned SQMR on $\mathbf{L}$ are shown in Fig. 6. The iterations are terminated when the relative residual goes below $10^{-8}$, and the legend indicates the type and number of smoothers used for the boundary and interior regions. We consistently apply a symmetric distributive Gauss-Seidel smoother only once in each V-cycle to the interior region. This approach significantly reduces the execution time per V-cycle, given that the boundary degrees of freedom are typically less than $3\%$ of the total degrees of freedom. Table. 1 provides more detailed information about the percentage of the number of boundary degrees of freedom.
The convergence plots reveal that our preconditioner achieves faster convergence compared to the pure multigrid method. Furthermore, the multiplicative Vanka smoother is more efficient than the addictive one, even when the latter is applied three times per V-cycle. The convergence rates of both multiplicative and addictive methods closely match the ideal scenario where a direct solver is employed for the boundary domains.
@@
6.1.2 Poiseuille Flow around Cylinder Example
In the case of 2D Poiseuille flow, it represents the steady, laminar flow of an incompressible fluid through a channel driven by a pressure gradient. This flow scenario is commonly studied in benchmarks schafer1996benchmark ; nicolas2011benchmark , particularly with the presence of a cylindrical obstacle. In our 2D scenario, the channel dimensions are $2.2m\times 0.41m$, and a cylindrical obstacle is placed at position $(0.2m,0.2m)$ with a radius of $0.05m$. The inflow boundary conditions are described by
@@ 6.1.2 Poiseuille Flow around Cylinder Example
where $\bar{u}=0.3m/s$ is the average horizontal velocity, and $H=0.41m$ is the width of the channel. On the outflow boundary, a homogeneous Neumann boundary condition is applied, while no-slippery boundary conditions are enforced on the top and bottom sides. The computational domain has dimensions $2200\times 410$, and a 6-level multigrid is used to solve the problem with $h=0.001m$. The partially solved velocity field and the convergence results are shown in Fig. 7, which shows that the effectiveness of symmetric multiplicative Vanka smoother. Besides that, the scalability of our multiplicative preconditioner and multigrid is demonstrated through various domain sizes in Fig. 8.
@@ 6.1.3 Poiseuille Flow around Hollow Square Example
In this example, we explore a more complex scenario involving Poiseuille flow around a hollow square obstacle. While pure multigrid suffices for simple applications, it may encounter divergence in more complicated scenarios. Here we use similar boundary conditions as those used in the cylindrical obstacle with $\bar{u}=1.0m/s$. The key difference lies in the presence of a hollow square obstacle, as depicted in Fig. 9 a). We perform our simulations on a $1024^{2}$ computational domain with an 8-level multigrid. It is observed that only the multiplicative preconditioner yields fast convergence, while the additive preconditioner and the pure multigrid method fail to converge. This behavior can be attributed to the close of the hollow square as the multigrid coarsens. Consequently, the solutions derived at coarser levels become increasingly inaccurate. However, when employed as a preconditioner, the multigrid method continues to guide SQMR towards successful convergence.
@@ 6.1.4 Brancher Example
In the hollow square example, we explored the limitations of pure multigrid methods. Now we present a more practical example to emphasize the defects of pure multigrid methods and the reliability of multiplicative preconditioner. We keep the same boundary conditions and computational domain as those used in the hollow square obstacle. Instead of a single obstacle, we apply multiple rectangular obstacles near the outflow boundary to branch a single inflow into several outflows, forming a structure referred to as a ”brancher”. As shown in 10 b), our multiplicative preconditioner converges very fast. In contrast, the addictive preconditioner and the pure multigrid methods still struggle to converge. The hollow square example and the brancher example clearly show the effectiveness and reliability of the multiplicative preconditioner. This fact leads us to focus on the multiplicative methods in our upcoming 3D examples, without considering the addictive method. Additionally, we investigate the impact of the numbers of boundary smoothers on the convergence rate of multiplicative preconditioner in 11. The result indicates that increasing the numbers of boundary smoother leads to faster convergence rates.
@@
In this subsection, we consider four 3D examples using only multiplicative preconditioner and multigrid. Driven cavity example and Poiseuille flow with a cylindrical obstacle example show that our multigrid preconditioner achieves comparable convergence rates to pure multigrid methods. Then the last two examples highlight the capability of our approach to address more complicated scenarios beyond the capabilities of pure multigrid solvers.
@@ 6.2.1 Driven Cavity Example
3D driven cavity example is similar to its 2D version. In this case, we consider a unit cube instead of a unit square as our computational domain, applying no-slip boundary conditions to all sides except the top. The top side is subject to a constant horizontal velocity of $u=1.0m/s$. A 5-level multiplicative multigrid is used within a $128^{3}$ unit cube. We present the sliced velocity field and the convergence results in Fig. 12. We can see that our preconditioner converges faster compared to the multigrid method, closely resembling where the boundary domains are solved using a direct solver.

### v3

@@ 3.2 Symmetric Vanka smoother
The Vanka smoother is more powerful and broadly applicable, but incurs a higher computational cost per iteration compared to the distributive smoother. It can be viewed as a specific type of Schwarz smoother [32]. The fundamental idea behind the Vanka smoother is to solve the Stokes equations locally, subdomain by subdomain, wherein all degrees of freedom within one subdomain are updated simultaneously. Like other smoothers, Vanka relaxation can be expressed in the format of Equation 3. In the Vanka smoother, we divide the entire discrete domain $D$ into $n$ potentially overlapping subdomains $\{D_{1},...,D_{n}\}$ as shown in Fig. 4. Each subdomain can be viewed as a local saddle-point problem. Solving each subdomain problem sequentially, akin to Gauss-Seidel, is known as multiplicative Vanka relaxation. The smoother operator can be defined as
@@
where $\mathbf{C}_{i}$ represents the selection matrix from domain $D$ to domain $D_{i}$. Alternatively, solving each subdomain problem simultaneously as Jacobi method, is known as the additive Vanka smoother. The smoother operator for this approach is defined as
@@
We prove that the additive Vanka smoother is always symmetric in B. To ensure the symmetry in the multiplicative Vanka smoother, we adopt a similar approach to symmetric Gauss-Seidel. We solve each subdomain from domain $D_{1}$ to $D_{n}$, and then in reverse order from $D_{n}$ to $D_{1}$ again. The proof of symmetry for our symmetric multiplicative Vanka smoother is given in C.
@@ 3.3 Symmetric integration of smoother
As previously discussed, these two smoothers provide advantages depending on the region of the computational domain. The symmetric distributive smoother provides reliable convergence and attractive computational cost when applied to the interior region, while the symmetric Vanka smoother exhibits greater suitability near the boundary, albeit at a higher cost. However, despite the fact that each can be made symmetric on its own, it’s not trivially true that any combination will inherently retain symmetry. To ensure strict symmetry with the multigrid framework, we propose the following smoothing step algorithm:
First, we partition the unknowns into two non-overlapping sets: boundary set $V_{1}$ and interior set $V_{2}$. Then we apply symmetric Vanka smoother to $V_{1}$, symmetric distributive smoother to $V_{2}$ and another symmetric Vanka smoother to $V_{1}$ again. The detailed algorithm is shown in Alg. 3, with a formal proof for its symmetry provided in D. In practice for our implementation, DGS smoother is used instead of DWJ, because it achieves a more attractive convergence rate but a similar runtime cost.
This approach leverages the strengths of both smoothers. On one hand, the larger size of $V_{2}$ ensures that the symmetric Vanka smoother isn’t overused, enhancing computational efficiency. On the other hand, employing symmetric Vanka smoother near the boundary helps prevent the potential divergence issues of the symmetric distributive smoother. Therefore, this integrated approach is a balanced and effective choice in practical applications compared to using either smoother alone.
@@
3:procedure Smooth($\mathbf{L}^{h}$, $\mathbf{x}^{h}_{0}$, $\mathbf{b}^{h}$)
@@ 4 Preconditioning
Although the use of the Vanka smoother significantly improves convergence near the boundary, multigrid is more stable when used as the preconditioner for a Krylov subspace solver. SQMR is known for its lower computational time and storage requirements, and supports arbitrary symmetric preconditioners, including indefinite ones from a multigrid V-cycle. Therefore, we propose using our multigrid as the preconditioner for SQMR. The application of the multigrid preconditioner is outlined in Alg. 4. It’s worth mentioning that not just any multigrid V-cycle can be used as the preconditioner. This procedure must be equivalent to $\mathbf{t}\leftarrow\mathbf{M}^{-1}\mathbf{s}_{0}$ where matrix $\mathbf{M}$ is symmetric. Fortunately, as we prove in the appendix, our multigrid formulation indeed satisfies this condition with the initial guess $\mathbf{x}_{0}=\mathbf{0}$.
@@
where $\gamma$ is an adequately small number. This modification corresponds to the classical penalty method. Since the modification is only applied as a preconditioner instead of as a replacement of the system matrix, the divergence-free property can be always exactly maintained.
@@
Alg. 4 Multigrid-Preconditioned SQMR
@@ 5 Discrete Domain Design
In this section, we describe our design decisions for the discrete description of the computational domains we target. We highlight that, as a matter of scope, we focus on a first-order accurate discretization of the Stokes equations, at grid-cell precision. This is primarily a choice for simplicity of exposition; our solver architecture and specifically our use of Vanka relaxation near the boundary could have been combined with a higher order scheme to provide sub-cell resolution of boundary conditions if desired. In addition, even if the finest level of discretization in the multigrid hierarchy was discretized at higher order, it would still be an option to use first-order accurate discretizations at coarser levels, especially when the multigrid cycle is intended as a preconditioner.
We adopt a domain description where every cell is labeled as Dirichlet, exterior or interior as illustrated in Fig. 5. Using this designation of cells as input, we proceed to also classify velocity or pressure variables as “Dirichlet”, “active” or ”inactive” as well. Specifically, any velocity variables on faces of Dirichlet cells are treated as Dirichlet boundary conditions. Velocities on faces between two “interior” cells are treated as active degrees of freedom and are solved by our scheme. Any velocity variables that do not neighbor an active cell are treated as inactive, and excluded from our solve. Finally, pressures at centers of interior cells are active degrees of freedom, while others (at centers of Dirichlet or exterior cells) are inactive and excluded from the solve.
@@
We adopt this cell-based designation, as opposed to, e.g. designating individual grid faces as Neumann faces, because our design leads itself naturally to coarsening this cell-level designation as follows: A cell at a coarser level inherits its label based on the labels of the finer-level cells within it. Specifically, a coarser-level cell is labeled as Dirichlet if it encloses any finer-level Dirichlet cell, otherwise as interior if it contains any finer-level interior cell. A cell remains an exterior one if all enclosed finer-level cells are exterior. This discrete domain design enables a straightforward partition of degrees of freedom into two non-overlapping sets for the use of symmetric Vanka and symmetric distributive smoothers. If a cell is proximal to any Dirichlet cell or exterior cell, all the degrees of freedom on this cell are smoothed by symmetric Vanka smoother. Conversely, symmetric distributive smoother is used for all other degrees of freedom.
Every interior velocity (on the face between two interior cells) or pressure variable is paired with a discrete equation, according to a finite difference scheme. Dirichlet velocities are similarly paired with a Dirichlet boundary condition. Velocity variables on a face between one interior and one exterior cell are treated as Neumann conditions; we treat those Neumann conditions as zero in our numerical examples, with the understanding that any nonzero conditions can be moved to the right-hand-side of the discretized equations without further alteration of the discrete operator.
6 Numerical Convergence Experiments
In this section, we present the results of our numerical convergence experiments, which compare the performance of our method with the pure multigrid method under several different boundary scenarios. We set the fluid viscosity to $\eta=10^{-3}m^{2}/s$, and $\gamma=10^{-3}$. To enhance the efficiency of Vanka smoothing, we pre-factorize $3^{4}$ in 2D or $3^{6}$ in 3D local saddle point problems, based on whether neighboring cells are Dirichlet, exterior or interior. It’s straightforward to parallelize the additive Vanka smoother, and an 8-color scheme is applied to parallelize the multiplicative Vanka smoother. All examples are implemented in C++ code, and we utilize Intel Pardiso and Eigen as the direct solvers for the bottom level of the V-cycle and for solving local saddle-point problems in the Vanka smoother respectively. The experiments are performed on a computer with an Intel(R) Xeon(R) CPU E5-2650 v3 @ 2.30GHz and 128 GB of main memory. Paraview is used for postprocessing the results. In the following section, we’ll discuss the results in 2D examples and 3D examples separately.
@@ 6 Numerical Convergence Experiments
In this subsection, we explore four 2D examples to assess the performance of our approach. Specifically, we examine the driven cavity example and the Poiseuille flow around a cylinder example to evaluate the convergence rates in comparison to pure multigrid methods. Additionally, the remaining two examples demonstrate the capability of our method to effectively handle more complex scenarios beyond the scope of pure multigrid methods.
@@ 6.1.1 Driven Cavity Example
The driven cavity example is a well-established benchmark in the field of computational fluid dynamics. In the 2D scenario, the computational domain is a unit square, where no-slip boundary conditions are imposed on the left, right and bottom sides. On the top side, a zero vertical velocity and a constant horizontal velocity of $u=1.0m/s$ is imposed. We use a $1024^{2}$ unit square as our computational domain and an 8-level multigrid as our preconditioner and solver. The solved velocity field with streamline and the convergence of pure multigrid on $\mathbf{\tilde{L}}$ and multigrid-preconditioned SQMR on $\mathbf{L}$ are shown in Fig. 6. The iterations are terminated when the relative residual goes below $10^{-8}$, and the legend indicates the type and number of smoothers used for the boundary and interior regions. We consistently apply a symmetric distributive Gauss-Seidel smoother only once in each V-cycle to the interior region. This approach significantly reduces the execution time per V-cycle, given that the boundary degrees of freedom are typically less than $3\%$ of the total degrees of freedom. Table. 1 provides more detailed information about the percentage of the number of boundary degrees of freedom.
@@
The convergence plots reveal that our preconditioner achieves faster convergence compared to the pure multigrid method. Furthermore, the multiplicative Vanka smoother is more efficient than the additive one, even when the latter is applied three times per V-cycle. The convergence rates of both multiplicative and additive methods closely match the ideal scenario where a direct solver is employed for the boundary domains.
6.1.2 Poiseuille Flow around A Cylinder Example
The case of 2D Poiseuille flow represents the steady, laminar flow of an incompressible fluid through a channel driven by a pressure gradient. This flow scenario is commonly studied in benchmarks [33, 34], particularly with the presence of a cylindrical obstacle. In our 2D scenario, the channel dimensions are $2.2m\times 0.41m$, and a cylindrical obstacle is placed at position $(0.2m,0.2m)$ with a radius of $0.05m$. The inflow boundary conditions are described by
@@ 6.1.2 Poiseuille Flow around A Cylinder Example
where $\bar{u}=0.3m/s$ is the average horizontal velocity, and $H=0.41m$ is the width of the channel. On the outflow boundary, a homogeneous Neumann boundary condition is applied, while no-slip boundary conditions are enforced on the top and bottom sides. The computational domain has dimensions $2200\times 410$, and a 6-level multigrid is used to solve the problem with $h=0.001m$. The partially solved velocity field and the convergence results are shown in Fig. 7, which shows that the effectiveness of the symmetric multiplicative Vanka smoother. Additionally, the scalability of our multiplicative preconditioner and multigrid is demonstrated through various domain sizes in Fig. 8.
@@ 6.1.3 Poiseuille Flow around Hollow Square Example
In this example, we explore a more complex scenario involving Poiseuille flow around a hollow square obstacle. While pure multigrid suffices for simple applications, it may encounter divergence in more complicated scenarios. Here we use similar boundary conditions as those used in the cylindrical obstacle with $\bar{u}=1.0m/s$. The key difference lies in the presence of a hollow square obstacle, as depicted in Fig. 9. We perform our simulations on a $1024^{2}$ computational domain with an 8-level multigrid. It is observed that only the multiplicative preconditioner yields fast convergence, while the additive preconditioner and the pure multigrid method fail to converge. This behavior can be attributed to the closing of the hollow square as the multigrid coarsens. Consequently, the solutions derived at coarser levels become increasingly inaccurate. However, when employed as a preconditioner, the multigrid method continues to guide SQMR towards successful convergence.
@@ 6.1.4 Brancher Example
In the hollow square example, we explored the limitations of pure multigrid methods. Now we present a more practical example to emphasize the defects of pure multigrid methods and the reliability of the multiplicative preconditioner. We keep the same boundary conditions and computational domain as those used in the hollow square obstacle. Instead of a single obstacle, we apply multiple rectangular obstacles near the outflow boundary to branch a single inflow into several outflows, forming a structure referred to as a “brancher”. As shown in Fig. 10, our multiplicative preconditioner converges very fast. In contrast, the additive preconditioner and the pure multigrid methods still struggle to converge. The hollow square example and the brancher example clearly show the effectiveness and reliability of the multiplicative preconditioner. This fact leads us to focus on the multiplicative methods in our upcoming 3D examples, without considering the additive method. Additionally, we investigate the impact of the number of boundary smoother iterations on the convergence rate of the multiplicative preconditioner in Fig. 11. The result indicates that increasing the numbers of boundary smoother iterations leads to faster convergence rates.
@@
In this subsection, we consider four 3D examples using only multiplicative preconditioner and multigrid. The driven cavity example and the Poiseuille flow with a cylindrical obstacle example show that our multigrid preconditioner achieves comparable convergence rates to pure multigrid methods. The last two examples highlight the capability of our approach to address more complicated scenarios beyond the capabilities of pure multigrid solvers.
@@ 6.2.1 Driven Cavity Example
The 3D driven cavity example is similar to the 2D version. In this case, we consider a unit cube instead of a unit square as our computational domain, applying no-slip boundary conditions to all sides except the top. The top side is subject to a constant horizontal velocity of $u=1.0m/s$. A 5-level multiplicative multigrid is used within a $128^{3}$ unit cube. We present the sliced velocity field and the convergence results in Fig. 12. We can see that our preconditioner method converges faster than the multigrid method, closely resembling where the boundary domains are solved using a direct solver.